OpenStudy (anonymous):

A recipe for lasgna calls for 93/125 lb of beef per person. Express this as a decimal. You want to use this recipe to make enough lasgna for 75 people. If beef at your grocery store costs $2 per pound, how much would you spend on the beef? "The recipe calls for ___ lb of beef per person. (Type an integeror a decimal.)

OpenStudy (igreen):

Okay, so you need 93/125 lb of beef per person. 93/125 = 0.744. So that's 0.744 lb of beef per person. If you want to make it for 75 people, you will multiply 75 * 0.744. 75 * 0.744 = 55.8 lb. So you need 55.8 lb of beef for 75 people. If beef at the grocery store costs $2 a pound, Multiply: 55.8 * 2 = $111.60 for 55.8 lb of beef. The term that goes in the blank will be '0.744' lb of beef per person.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Niners49

OpenStudy (anonymous):

Thank you. I was completely stumped on this question. But that makes perfect sense now. Thanks again.